# On the Retention of Edited Knowledge in Fine-Tuned Language Models

### v1

@@ B.2 Hyperparameter for finetuning
Unstructured finetuning: For GPT-2 XL, in align with its pretraining stage, we employ batch size of 256, AdamW optimizer with learning rate of 5e-5, beta of (0.9,0.999) and weight decay of 0.01. The training data set contains 60k data, and the validation data set contains 1k data. The model is evaluated on the validation set every 20 steps. The training stops once the validation loss goes below 3.1.
@@
Structured finetuning: We employ batch size of 256, AdamW optimizer with learning rate of 2e-5, beta of (0.9,0.999) and weight decay of 0.01. The training data set contains 3,000 data, and the validation data set contains 100 data. The model is evaluated on the validation set every 20 steps. The training stops once the accuracy on the validation set goes above 70%.
@@
Supervised finetuning(SFT): For GPT-2 XL, we use an AdamW optimizer with learning rate of 2e-5, with beta of (0.9,0.999) and weight decay of 0.01. The training data set contains 10,000 data, and the validation data set contains 1,000 data. Training is performed with a batch size of 64, and the model is evaluated on the validation dataset every 20 steps. We halt training once the validation loss falls below 1.7.
@@
Appendix C Experiment results of Llama3-8B
Edit method ROME FT MALMEN
Edited knowledge 0.374 0.634 0.815
Intrinsic knowledge 0.824 0.831 0.817
Appendix D Experiment results for different edit layer

### v2

@@ C.3 Hyperparameter for fine-tuning
Unstructured fine-tuning: For GPT-2 XL, in align with its pre-training stage, we employ batch size of 256, AdamW optimizer with learning rate of 5e-5, beta of (0.9,0.999) and weight decay of 0.01. The training data set contains 60k data, and the validation data set contains 1k data. The model is evaluated on the validation set every 20 steps. The training stops once the validation loss goes below 3.1.
@@
Structured fine-tuning: We employ batch size of 256, AdamW optimizer with learning rate of 2e-5, beta of (0.9,0.999) and weight decay of 0.01. The training data set contains 3,000 data, and the validation data set contains 100 data. The model is evaluated on the validation set every 20 steps. The training stops once the accuracy on the validation set goes above 70%.
@@
Supervised fine-tuning(SFT): For GPT-2 XL, we use an AdamW optimizer with learning rate of 2e-5, with beta of (0.9,0.999) and weight decay of 0.01. The training data set contains 10,000 data, and the validation data set contains 1,000 data. Training is performed with a batch size of 64, and the model is evaluated on the validation dataset every 20 steps. We halt training once the validation loss falls below 1.7.
@@
C.4 Knowledge Retention Rate
The model can sometimes assigns the highest probability to a stop word but not the target word. However, this does not necessarily imply a lack of knowledge. In such cases, we employ greedy decoding to generate subsequent tokens until a non-stop token is produced.
C.5 Experiment Results for Different Edit Layer